OpenStudy (anonymous):

Convert the map scale to a unit rate. Show your work. Interpret the meaning of the unit rate 3/10 in.= 7/8 mi

OpenStudy (ranga):

3/10 in.= 7/8 mi Convert to unit rate means 1 in = how many miles? Multiply both sides by 10 and divide both sides by 3.

OpenStudy (ranga):

After you calculate 1 in = how many miles (call it x), the interpretation is: If I measure 1 inch between two points on the map then in real life measurement it will be x number of miles between those two points.

OpenStudy (anonymous):

When I multiplied bother sides I got 30/100=70/80 what do I do now?

OpenStudy (ranga):

The left side is 3/10. What is 3/10 * 10 ?

OpenStudy (anonymous):

I got 30/100

OpenStudy (ranga):

You don't multiply both the numerator and the denominator. 3/10 is a fraction. When you multiply by 10, the 10 in the top and the ten in the bottom will cancel out.

OpenStudy (anonymous):

Oh

OpenStudy (ranga):

3/10 in = 7/8 mi multiply both sides by 10 3/10 * 10 in = 7/8 * 10 mi 3 in = 70/8 mi divide both sides by 3 1 in = 70/8 / 3 = 70/24 = 35/12 mi So the unit rate is: 1 inch = 35/12 miles

OpenStudy (ranga):

The whole reason behind multiplying by 10 and dividing by 3 is to make the left hand side 1 inch (by getting rid of the fraction 3/10)

OpenStudy (ranga):

The interpretation of the unit rate 1 inch = 35/12 miles is that for every inch measured on the map it represents 35/12 miles in real life. If the distance on the map between two points is i inch then in real life the distance between those two points will be 35/12 miles.

OpenStudy (anonymous):

Thanks I will give you a medAl tomorrow

OpenStudy (ranga):

you are welcome.